Divers know that the pressure exerted by the water above them increases about 100kPa with every 10.2m of depth. This means that at 10.2 m below the surface, the pressure is 201 kPa, at 20.4 m below the surface it’s 301 kPa & so forth. Given that the volume of a balloon is 3.5 L at STANDARD TEMPERATURE & PRESSURE (273 K & 1 atm), what is the volume of the balloon 51 m below the water’s surface if the temperature remains constant?
1
Expert's answer
2020-10-22T08:37:00-0400

pV=nRT.

n=pV/RT = 101.325 x 3.5 / 8.314 x 273= 0,156247109 mol.

51 x 100/10.2 = 500 kPa.

V=nRT/p = 0,156247109 x 8.314 x 273 / 500 = 0,709275 L.